An hydrochloric solution of unknown concentration is titrated using 0.150 M sodium hydroxide solution. If it took 54.3 mL of the sodium hydroxide solution to titrate a 25.0 mL sample of the hydrochloric acid solution, what is the concentration of the hydrochloric solution?
1
Expert's answer
2020-05-26T14:20:57-0400

The equation is HCl + NaOH = NaCl + H2O

c(NaOH)*V(NaOH) = c(HCl)*V(HCl)

c(HCl) = c(NaOH)*V(NaOH)/V(HCl) = (0.150 mol/L*54.3 mL)/25.0 mL = 0.326 mol/L